OpenStudy (konradzuse):

derivative of 4/(e^x +e^-x)

OpenStudy (konradzuse):

\[\frac{4}{e^x+e^{-x}|}\]

OpenStudy (konradzuse):

the quotient rule should be vu'-uv'/v^2

OpenStudy (anonymous):

Start with\[4\frac{d}{dx}(\frac{1}{e^x})+4\frac{d}{dx}(\frac{1}{e^{-x}})\] Use the quotient rule to solve these two derivatives

OpenStudy (konradzuse):

4e^(-x) would be -4e^(-x) and 4e^x would be the same.

OpenStudy (cwrw238):

sorry - don't know what i was thinking of then - you can use the quotient rule as u said

OpenStudy (konradzuse):

the quotient rule would be (e^x+ e^-x) *0(or is it just null) - 4(e^x - e^-x) all over (e^x+e^-x)^2

OpenStudy (konradzuse):

-4(e^x-e^-x) ------------- (e^x + e^-x)^2

OpenStudy (cwrw238):

v*du = 0 yes u dv = 4 * (e^x + (-e^(-x) = 4(e^x - e^(-x)) - thats the correct numerator i think

OpenStudy (cwrw238):

ah - its - 4(e^x - e(^-x)) - u r right

OpenStudy (anonymous):

While the quotient rule is the most formal way of calculating the derivatives of inverse exponential functions, you can do it more quickly by remembering this: \[\frac{d}{dx}(e^{ax})=ae^{ax}\] For example:\[\frac{d}{dx}(e^{-17x})=-17e^{-17x}\]